The length of a rectangle is three and one half times the width. If the area of the rectangle is 12m squared, find the width. I already have the answer from the key but would really like to know how to do it correctly. Thank you
Answer by mananth(16946)

width = x
length = 3.5 x
Area = 12 m^2
Area = L*W
12=x*3.5x
12=3.5x^2
3.4285=x^2
x= 1.85
width 1.85
length = 3x = 6.47